### v1

# Expert System for Bitcoin Forecasting: Integrating Global Liquidity via TimeXer Transformers

Abstract
Bitcoin price forecasting is characterized by extreme volatility and non-stationarity, making it hard for univariate time-series prediction models over long horizons. By integrating Global M2 Liquidity—aggregated from 18 major economies as exogenous variable we improve the predictions over long horizons.
Using TimeXer, we compare a liquidity-conditioned univariate forecasting model (TimeXer-Exog) against state-of-the-art univariate models, LSTM, N-BEATS, PatchTST, and standard TimeXer. Experiments conducted on daily Bitcoin price data from January 2020 to August 2025 demonstrate that explicit macroeconomic conditioning significantly stabilizes long-horizon forecasts.
At a 70-day forecast horizon, the proposed TimeXer-Exog model achieves a Mean Squared Error (MSE) of $10.814\times 10^{7}$ which outperformed the univariate TimeXer baseline by over 89%. The results show that explicitly conditioning deep learning models on Global Liquidity as an exogenous variable gives substantial improvements in long-horizon Bitcoin price forecasting.
Keywords: Bitcoin forecasting, Global liquidity, TimeXer, Transformers, Deep learning
1 Introduction
Early attempts to forecast Bitcoin prices were done using statistical tools, especially ARIMA models (Sagheer2025; Xu2025). ARIMA is good at picking up short-term linear patterns, like the ones in a relatively calm equity index over a few weeks. However, bitcoin does not behave like that. Its price is noisy and very volatile, which statistical models and tools fail to capture. ARIMA requires stationarity and linear structure, this is exactly what bitcoin isn’t. Large rallies, sudden crashes, or regime shifts often get smoothed away (Pratas2023; Mizdrakovic2024). As a result, these models can perform well over a short horizon, their accuracy drops once forecast horizon is about 60 or 90 days (Mousa2025; Kareem2025).
Researchers moved RNN models, which were designed to handle sequences. LSTM and GRU models, were successful in learning the non-linear patterns from the data (Bagheri2025; Xu2025). They can, for example, react differently to a slow upward trend than to a sudden spike triggered by a regulatory headline. These models are not perfect. Because they process data step by step, they scale poorly and struggle when dependencies stretch far back in time. Capturing relationships across hundreds of days still remains hard, and unreliable (Zhao2022; Zheng2025).
More recently, attention has shifted toward Transformer-based models, time series forecasting moved to self-attention rather than strict sequence processing (Nie2022; Zhao2022). Instead of remembering the past one step at a time, these models can look across an entire window at once. PatchTST takes this idea further by breaking the time series into patches, small segments that preserve local structure while keeping computation manageable (Nie2022). For time series forecasting which require long look-back windows, Transformer models have repeatedly outperformed LSTM-based approaches, especially for long-horizon predictions (Nie2022; Zheng2025).
Long-term forecasting improves when models are allowed to look beyond price history itself. The TimeXer framework reflects this shift. It combines patch-wise attention for the main series with cross-attention for external variables, making it possible to integrate data that arrive at different frequencies and are not neatly aligned in time (Wang2024). This is closer to how real financial systems operate. Macroeconomic indicators do not update daily, yet markets respond to them gradually and sometimes unevenly.
Global liquidity, often measured by M2 money supply, has attracted particular attention (Gu2025). The logic is intuitive, even if the relationship is not perfectly stable. Empirical work using time-varying Granger causality suggests that the influence of M2 on Bitcoin prices comes and goes, strengthening during expansionary regimes and fading at other times (Gu2025). Cointegration analyses go a step further, pointing to a long-run equilibrium relationship in which Bitcoin prices respond more than proportionally to liquidity growth (Kokabian2025). Importantly, these effects do not appear overnight. The impact of M2 unfolds with lags, which fits the idea that liquidity works its way through portfolios and risk appetite before reaching crypto markets (Kokabian2025).
We wanted a forecasting approach that combines long-horizon modeling with explicit macroeconomic context. In this paper we integrate global M2 money supply as an exogenous variable within a Transformer-based framework. Using TimeXer and global liquidity as exogenous variable, we can have better bitcoin price predictions over long horizons.
From a broader financial engineering perspective, this direction reflects a quiet change in how intelligent systems are defined. Accuracy matters, but so does context. Analysts look at bitcoin prices often in relation to liquidity cycles, central bank policy, and broader risk sentiment. This is what differentiates a macroeconomist from a black-box predictor. Moving from a univariate setup to a multivariate expert framework, and by allowing the model to explicitly attend to global monetary conditions, the proposed approach aims to deliver insights that are more robust to the drift and structural change that plague many traditional forecasting tools.
2 Related Work
2.1 Bitcoin forecasting using ARIMA
Bitcoin price predictions was initially done using ARIMA, partly because it was already a familiar tool in financial time series analysis (Kareem2025; Xu2025). For short-term movements, especially when prices drift without too much drama, ARIMA can be very reliable. One can imagine fitting it over a calm stretch of daily log returns and getting something that looks sensible for the next few days. The trouble starts when the market becomes volatile. Bitcoin prices are rarely stationary for too long, sharp jumps or sudden collapses violate the linear structure ARIMA assumes (Sagheer2025; Pratas2023). During volatile episodes, the model has a habit of smoothing away extreme moves, as if the crash or rally were just noise rather than the main event (Kareem2025). Forecasts also lose credibility quickly as the forecast horizon increases, which makes ARIMA a weak candidate for longer-term planning (Mousa2025).
2.2 Bitcoin forecasting using CNN, RNN, and LSTM models
As attention shifted toward more flexible approaches, deep learning models began to dominate the discussion (Bagheri2025). RNNs, and LSTMs in particular, seemed well suited to the task. They were designed to remember patterns over longer periods of time, and in practice they handled Bitcoin’s nonlinear behavior better than classical models (Hochreiter1997; Mardjo2024). Some studies paired CNNs with LSTMs, letting convolutional layers pick up local price patterns, such as short bursts of momentum, before passing them to a recurrent structure (Sagheer2025; Bagheri2025). Interestingly, several comparative analyses found that GRUs often matched or even outperformed LSTMs, while using fewer parameters and training more efficiently (Kaur2025; Mohammadjafari2024; Nafisah2025). These models share a common weakness. Because they process sequences step by step, they become cumbersome when dependencies stretch far back in time, which is often the case in crypto markets which are shaped by long liquidity cycles.
2.3 Bitcoin forecasting using Transformer-based models
Transformer models entered the scene as a response to these limitations (Zheng2025). By relying on self-attention rather than strict sequential processing, they can examine long windows of data in parallel. This makes it easier to capture distant dependencies without the training bottlenecks typical of RNNs. In volatile settings like cryptocurrency markets, time-series Transformers have delivered encouraging results (Zheng2025). In PatchTST we divide the input series into patches, a practical compromise that preserves local structure while keeping attention costs under control (Nie2022). Across a range of long-horizon tasks, transformer models have repeatedly outperformed LSTM baselines, producing forecasts that are not only more accurate but also less erratic over time (Nie2022). That said, architecture alone does not explain everything, and context still matters.
2.4 Incorporating exogenous variables in Bitcoin forecasting
Long-term Bitcoin forecasting rarely works if only price history is considered. Researchers have experimented with on-chain activity, sentiment measures, and macroeconomic indicators to reflect broader market forces (Mahfooz2024; Grubisic2025). Among these, global liquidity, typically proxied by M2 money supply, stands out (Gu2025). The connection is not mechanical, but there is evidence of a long-run equilibrium relationship between M2 and Bitcoin prices, with the influence becoming more pronounced during periods of monetary expansion (Gu2025; Kokabian2025). Importantly, these effects tend to emerge with delays rather than instant reactions. Transformer based models like TimeXer make it easier to incorporate such variables by allowing cross-attention between price dynamics and external drivers (Wang2024). This line of work points toward forecasting systems that treat Bitcoin not as an isolated series, but as part of a wider monetary environment, where global liquidity an play a role in shaping long-horizon price behavior.
3 Methodology
3.1 The TimeXer Architecture
TimeXer was built with a fairly practical goal in mind: making it easier to bring external information into time series forecasting without tearing apart the Transformer framework. TimeXer reuses the familiar Transformer backbone and adapts how inputs are represented and attended to. The result feels less like a radical redesign and more like a careful rethinking of how different data streams should talk to each other.
Figure 1: The TimeXer architecture showcasing the integration of endogenous and exogenous data streams through specialized embedding and attention mechanisms.
Exogenous Input Integration and Cross-Attention.
TimeXer is composed of three token types to capture both internal temporal dynamics and external causal factors: patch-level temporal tokens and a learnable global token for the endogenous Bitcoin series, and series-wise variate tokens for exogenous variables.
  1. 1.
    Endogenous: The primary Bitcoin price series (endogenous variable) is split into non-overlapping patches to finely capture temporal dependencies. The learnable global token for the endogenous series, acts as a macroscopic representation to aggregate information across the entire series.
  2. 2.
    Exogenous: Global Liquidity is embedded as series-wise variate tokens. This representation handles practical irregularities which exist in real-world data, such as temporal misalignment, missing values, and mismatched frequencies.
As shown in Figure 1, the model processes the input through two parallel pathways before combining them in the Transformer backbone. The endogenous (a) divides the Bitcoin price series into non-overlapping patches and adds a randomly initialized global token, forming the main temporal representation. At the same time, the exogenous (b) processes the Global Liquidity series separately, embedding it into variate-wise tokens that retain information about the external macroeconomic driver.
Both representations are then passed to the Transformer backbone. Here, patch-wise self-attention (c) captures temporal patterns within the Bitcoin series, while variate-wise cross-attention (d) uses the endogenous global token to selectively incorporate information from the exogenous variables. This design allows the model to learn Bitcoin’s internal dynamics while continuously adjusting to the broader macroeconomic signals from global liquidity.
3.2 Baseline Models
To benchmark performance, several established deep learning were used for univariate forecasting:
  • TimeXer (Endogenous Only): This configuration serves as a control, testing the framework’s native capability by relying solely on temporal patterns extracted from the Bitcoin price series.
  • PatchTST: This Transformer variant utilizes subseries-level patches along with channel independence to enhance long-term forecasting accuracy (Nie2022).
  • LSTM: A foundational recurrent architecture designed to overcome the vanishing gradient problem and efficiently capture long-term dependencies in sequential data (Hochreiter1997; Mahfooz2024).
  • N-BEATS: A purely deep learning architecture consisting of a very deep stack of fully-connected layers linked by residual connections (Oreshkin2019).
Classical linear models such as ARIMAX are known to suffer from stationary constraints and recursive forecast instability in highly volatile and non-stationary financial series; they are included only as diagnostic baselines and are not expected to compete with attention-based models at extended forecast horizons.(melina2024comparative)
3.3 Bitcoin and Global Liquidity Relationship
The integration of Global M2 Liquidity is justified by its documented long-run influence on Bitcoin pricing. A cointegration analysis established a strong relationship, characterized by a long-run elasticity estimate of 2.65, meaning a 1% increase in the M2 money supply is associated with a 2.65% increase in the price of Bitcoin (Kokabian2025). The stability of this long-run relationship is validated by a statistically significant Error Correction Term (ECT), $\lambda^{\prime}=-0.12$ (Kokabian2025). The magnitude of this coefficient demonstrates that approximately 12% of any deviation from the long-run equilibrium is corrected monthly.
Figure 2: Bitcoin open price v/s Global Liquidity with a 12-week lead (both are standard normalized)
Figure 2 highlights a clear lead-lag relationship between Global Liquidity and Bitcoin’s price action. Shifting the M2 liquidity data back by 12 weeks, the alignment of major cyclical peaks and troughs can clearly be observed. This observation suggests that changes in the global liquidity acts as a leading indicator, with a transmission delay of approximately one fiscal quarter before the liquidity is fully reflected in Bitcoin’s price. TimeXer leverages this 12-week lead as a predictive ”look-ahead” window, thus improving the predictions.
3.4 Statistical Evaluation: Model Confidence Set (MCS)
To evaluate the predictive performance of the proposed TimeXer-Exog model against the set of benchmark models ($\mathcal{M}_{0}=\{\text{TimeXer, PatchTST, LSTM, N-BEATS}\}$), we use Model Confidence Set (MCS) Hansen2011MCS.
Unlike traditional tests (e.g., the Diebold-Mariano test (Diebold1995)), which can suffer from the multiple comparisons problem and data snooping bias when many models are evaluated (White2000; Sullivan1999), the Model Confidence Set (MCS) procedure determines a subset of models $\mathcal{M}^{*}\subset\mathcal{M}_{0}$ that contains the ”true best” models with a specified confidence level $(1-\alpha)$. We iteratively eliminate models found to be significantly inferior until the null hypothesis of Equal Predictive Ability (EPA) cannot be rejected for the remaining set (Hansen2011MCS).
Since we use long horizon forecasting (Horizon $h=70$), the forecast errors are expected to exhibit serial correlation up to lag h1h-1. To account for this dependency structure and prevent spurious significance, we utilize the Block Bootstrap method (Kunsch1989) within the MCS framework. We set the block size equal to the forecast horizon ($h$) to preserve the internal correlation structure of the error series during resampling (Politis2004). We set the significance level $\alpha=0.10$. This means models with a p-value <0.10<0.10 are excluded from the Superior Set Model (SSM).
4 Data and Experimental Setup
4.1 Data Description
The endogenous dataset consists of daily Bitcoin (BTC) closing prices spanning from January 1, 2020, to January 21, 2025. This period has multiple market regimes, including monetary expansion phases, heightened volatility episodes, and prolonged drawdowns. This provides the perfect environment for evaluating long-horizon forecasting performance.
To ensure strict out-of-sample evaluation, the final 120 daily observations, from January 22, 2025, to May 21, 2025, were reserved exclusively for testing. All models were trained only on historical data prior to this window.
Experimental reference date and temporal alignment.
27 August 2025 was fixed as the reference date for this study, which represents the latest point in time at which both Bitcoin price data and Global Liquidity information were assumed to be fully observable. To prevent any form of look-ahead bias, all exogenous inputs were strictly restricted to values available on or before this date.
Since Global Liquidity was incorporated using a dense lag structure spanning 1 to 98 days, the effective sample end date is determined by the maximum lag applied. Specifically, applying a 98-day lag implies that the most recent usable Global Liquidity observation corresponds to 21 May 2025 (i.e., 98 days prior to 27 August 2025). Consequently, after lag construction and temporal alignment, the final row of the modeling dataset ends on 21 May 2025, even though the experiment itself is conceptually anchored to late August 2025.
This design ensures strict temporal causality: at every training and evaluation step, the model only has access to information that would have been available at the decision time, thereby eliminating forward-looking bias while preserving the full liquidity transmission window.
The exogenous variable employed in this study is Global Liquidity, constructed as an aggregate measure of broad money supply across a diversified set of major and emerging economies. Specifically, the indicator includes M2 (or the closest available broad money proxy) from the following economies: the United States, Euro Area, China, Japan, United Kingdom, Canada, Australia, India, South Korea, Brazil, Russia, Switzerland, Mexico, Indonesia, Turkey, Saudi Arabia, Argentina, and South Africa.
All national money supply series were converted into U.S. dollars using corresponding daily foreign exchange rates prior to aggregation. The resulting Global Liquidity measure therefore reflects worldwide monetary expansion in a unified currency framework, rather than being dominated by any single economy.
Empirical evidence suggests that changes in global liquidity affect Bitcoin prices with a delayed transmission. To account for this effect, the aggregated Global Liquidity series was shifted forward by 12 weeks, reflecting the estimated average market adjustment period through which liquidity conditions propagate into digital asset valuations.
4.2 Exogenous Input Construction
Although the 12-week shift captures the dominant lead–lag relationship between Global Liquidity and Bitcoin prices, the forecasting framework does not rely on a single fixed lag. Importantly, the forecast horizon $H$ denotes the length of the prediction window and does not correspond to the liquidity lag itself.
Rather than imposing a predetermined lag structure, a dense lag configuration was constructed to allow the model to learn the full liquidity transmission profile endogenously. Specifically, multiple lagged versions of the Global Liquidity series were generated at the following offsets:
$1,7,14,21,28,35,42,49,56,63,70,77,84,91,\text{ and }98\text{ days}.$
These lagged features are denoted as:
$\text{global\_lag}_{1},\text{global\_lag}_{7},\ldots,\text{global\_lag}_{98}.$
This dense lag structure enables the TimeXer architecture to infer both short- and long-delay effects of global liquidity on Bitcoin prices without explicitly specifying a functional form. Through variate-wise cross-attention, the model dynamically assigns importance to different liquidity lags depending on the forecast horizon, allowing it to adaptively capture regime-dependent transmission dynamics.
4.3 Training Setup
No manual normalization or scaling was applied during preprocessing. Instead, standardization was handled internally by each model through built-in normalization mechanisms such as Reversible Instance Normalization (RevIN) (Kim2022) used in PatchTST. This approach ensures that normalization statistics are computed independently for each sample, thereby eliminating potential information leakage across time.
All models were trained using the Mean Squared Error (MSE) loss function:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2},$
where $y_{i}$ and $\hat{y}_{i}$ denote the observed and predicted Bitcoin prices, respectively. MSE was selected because it penalizes large forecast errors heavily, which is particularly important in cryptocurrency markets which are filled with price swings.
4.4 Implementation Using the NeuralForecast Library.
All deep learning models were implemented using the NeuralForecast library, which provides a state-of-the-art framework for reproducing and training deep learning forecasting models (Olivares2022). The NeuralForecast framework was used to train and evaluate all the models (TimeXer, PatchTST, LSTM, and N-BEATS).
The library has native support for future exogenous variables, thus we could easily integrate lagged Global Liquidity features into the TimeXer-Exog model. Model training was conducted using horizon-specific fitting with early stopping based on validation loss Rolling-origin forecasting was employed to assess true out-of-sample performance. This standardized implementation ensured fair comparison across all the models and minimized confounding effects arising from differences in training or evaluation procedures.
4.5 Evaluation Metrics and Forecast Horizons
Forecasts were generated for fixed horizons of
$H=\{7,14,21,28,35,42,49,56,63,70\}\text{ days}.$
Performance was evaluated on the held-out out-of-sample test window using Mean Squared Error (MSE) as the evaluation metric. For numerical clarity and comparability across horizons, all the MSE values are scaled by a factor of $10^{7}$, with lower values indicating better predictions.
This evaluation method focuses on long-term stability and consistent price direction, rather than trying to match short-term fluctuations. This matches the study’s goal of understanding how macro-financial factors like global liquidity help improve long-term Bitcoin price predictions.
5 Results and Discussion
5.1 Quantitative Performance
Table 1: Comparative MSE ($\times 10^{7}$) Across Forecast Horizons (Lower is Better). Best performance per horizon is shown in bold.
Horizon (H) TimeXer-Exog TimeXer LSTM NBEATS PatchTST
70 10.814 101.952 29.120 77.276 147.566
63 12.363 85.436 318.285 77.960 128.121
56 5.363 62.699 26.653 23.399 45.092
49 6.912 20.000 12.100 14.263 33.567
35 4.580 9.447 22.720 17.933 22.971
28 5.588 7.333 9.060 16.923 15.374
21 4.087 3.585 26.804 14.579 13.263
14 4.288 5.395 21.284 12.181 5.781
7 3.160 4.351 2.810 4.843 2.028
For the longest horizon, 70 days, the difference is clear. TimeXer-Exog outperforms the univariate TimeXer baseline by more than 89 percent.
Looking at Table 1, a clear pattern emerges as the forecast horizon increases. In short windows, 7 to 14 days, the models that rely only on price history perform well. PatchTST outperforms TimeXer-Exog. Over one or two weeks, Bitcoin prices are mostly influenced by recent trends and short-term trading activity that are already reflected in the price data.
TimeXer-Exog begin to dominate after the one month mark. From 28 days onward, TimeXer-Exog has the lowest MSE and has great stability. There is less drifts and there seems to be a clearer sense of direction. Univariate models such as LSTM and NBEATS, performed poorly as the forecast horizon increased. Without outside information, their predictions either become flat or drift aimlessly, therefore showing the limits of using only price data to forecast.
The strong performance of TimeXer-Exog at longer horizons shows macro-financial is improving with the predictions. Providing lagged global M2 liquidity gives the model access to slow-moving forces that tend to lead market behavior rather than follow it. Using cross-attention, liquidity signals guide the prediction when price signals weaken, which seems to be the main reason for the improved predictions in long forecast horizons.
5.2 Qualitative Forecast Analysis
Figures 3, 4, and 5 offer a useful visual comparison between the model predictions, showing typical behavior at short, medium, and longer horizons.
At the shortest horizon ($H=7$), PatchTST tracks Bitcoin price most closely, as seen in Figure 3. This outcome feels intuitive. Over a single week, price dynamics are largely shaped by very local patterns, recent momentum, and short-lived market reactions. In this setting, macroeconomic signals move slowly in the background and add little immediate guidance. As a result, the liquidity-aware TimeXer-Exog model does not yet separate itself from purely price-driven approaches, and any advantage from exogenous information doesn’t help with the predictions.
Figure 3: Forecast comparison at horizon $H=7$.
At the medium horizon ($H=28$), the value of conditioning on macro-financial information becomes easier to see. In Figure 4, the TimeXer-Exog forecast stays close to the actual Bitcoin price path, picking up both the sharp drawdown in March and the rebound that followed into May. Other models relying only on past prices tend to react more slowly, smoothing over the decline and lagging the recovery. The result is a loss of directional accuracy and a forecast that feels slightly out of sync with the actual price of bitcoin.
Figure 4: Forecast comparison at horizon $H=28$.
By the time the horizon reaches $H=63$, the weaknesses of purely univariate models are obvious. In Figure 5, many of the baseline forecasts begin to wander, drifting away from the realized price path and losing any clear sense of direction. Some flatten out, others veer off entirely, and in both cases the connection to the underlying market dynamics weakens. The TimeXer-Exog model behaves differently, its forecasts stay aligned with the broader trend, preserving directional consistency and a coherent structure over time. This shows the significance of the role of lagged global liquidity. Because global liquidity changes slowly but gives early signals, it helps long-term forecasts when price data alone is no longer sufficient.
Figure 5: Forecast comparison at horizon $H=63$.
5.3 Statistical Significance and Robustness Analysis
Mean squared error is an easy way to rank models from best to worst. Still, it only scratches the surface. Prices jump for reasons that have little to do with forecasting skill, and a tiny edge in error can just as easily be luck as insight. That uncertainty tends to get glossed over when models are ranked mechanically.
This is where the Model Confidence Set becomes useful. Table 2 reports the corresponding p-values across different forecast horizons, which forces a pause before declaring winners. Within this setup, a p-value of 1.000 means a model survives every comparison and stays in the set of top performers. Lower values tell a less flattering story. Once the threshold drops below 0.10, the supposed edge starts to look fragile, the kind that might disappear if the sample were reshuffled or extended by a few months. At that point, exclusion is hard to avoid.
Seen this way, the procedure does not magically identify the “true” best model. Rather, it reins in overconfidence. It helps distinguish improvements that persist under repeated scrutiny from those that shine briefly, perhaps because a particular market phase happened to line up in their favor. That distinction matters, especially when forecasting models are expected to hold up beyond the specific slice of data that first made them look good.
Table 2: Model Confidence Set (MCS) P-Values across Forecasting Horizons. Bold values indicate inclusion in the 90% confidence set ($\alpha=0.10$). Models with $p<0.10$ are statistically significantly inferior.
Horizon ($h$) TimeXer-Exog (Ours) TimeXer PatchTST LSTM N-BEATS
7 0.288 0.092 1.000 0.288 0.092
14 1.000 0.671 0.671 0.186 0.186
21 0.694 1.000 0.065 0.202 0.204
28 1.000 0.305 0.077 0.305 0.108
35 1.000 0.075 0.009 0.055 0.000
42 1.000 0.202 0.001 0.443 0.443
49 1.000 0.071 0.063 0.311 0.063
56 1.000 0.086 0.086 0.086 0.033
63 1.000 0.004 0.000 0.000 0.003
70 1.000 0.005 0.000 0.037 0.000
Short-Term Dynamics ($h\in[7,21]$): Over shorter horizons, the MCS results point to a fairly crowded field. At $h=14$ and $h=21$, several models remain statistically indistinguishable, with the confidence set including TimeXer-Exog, TimeXer, LSTM. In practical terms, this suggests that when looking only a few weeks ahead, Bitcoin prices still carry a lot of their own momentum. Recent price history does much of the work. Well-designed univariate models, such as PatchTST, LSTM and NBEATS, can extract enough structure from autocorrelation alone. In these windows, adding global liquidity does not help much.
Long-Term Superiority ($h\in[56,70]$): Things change as the horizon increases. For forecasts at $h=63$ and $h=70$, the results in Table 2 are unambiguous. TimeXer-Exog is the only one in the confidence set, with a p-value of 1.000. All other models, including the base TimeXer and PatchTST, fall out well before the 10 percent threshold. At this point, the advantage of incorporating exogenous information is clearly evident.
Discussion of Results: This separation at longer horizons speaks directly to the role of global M2 liquidity. As forecasts horizons increase, the usefulness of past price movements fades. Autoregressive signals weaken, and univariate models begin to drift, often collapsing toward their own long-run averages or accumulating error along the way. The TimeXer-Exog model, drawing on a 12-week lagged liquidity signal, it retains access to information that is not present in recent prices. Because this signal leads the market rather than reacting to it, it offers a more stable reference point when price-based cues run thin. The MCS results suggest that the improvement in long-horizon accuracy is structural rather than accidental. In other words, the gains appear to come from how macroeconomic information is built into the model, not from statistical luck.
Data Availability
All data used here are publicly accessible. The Bitcoin prices, along with the processed global liquidity series, can be found in the gihub repository at https://github.com/sravankarthik/bitcoin-timeXer/blob/main/bitcoin.csv . The same repository also contains the full notebook needed to replicate the analysis, including the steps used to construct the global liquidity measure, introduce lagged features, and train all the models. The TradingView Pine Script for bitcoin and global liquidity is included as well https://github.com/sravankarthik/bitcoin-timeXer . With these materials, the experiments reported in this paper should be reproducible in full, or at least close enough.
6 Conclusion
The results point to a fairly clear takeaway. Using global M2 liquidity as exogeneous variable in TimeXer helps turn what might otherwise be a prediction model into something closer to an expert system for digital asset management. Over longer horizons, up to about 70 days, anchoring forecasts to a macroeconomic signal appears to matter a great deal. The observed reduction in forecast error, close to 89 percent relative to standard benchmarks, suggests that the model is not simply fitting price patterns more tightly, but is aligning its expectations with broader monetary conditions.
These results are beyond just academic accuracy metrics. This framework could mitigate risk in risk management systems or exist in algorithmic trading pipelines. Using global liquidity we can prevent the model from overreacting to short-lived price shocks, the sort of sudden moves that often dominate crypto markets for a short time and then fade.
Long-horizon forecasting in crypto markets or any market seems unlikely to work if it relies only on past prices. Past prices are important, but without explicit links to macroeconomic drivers, forecasts tend to drift. Attention-based architectures that incorporate exogenous liquidity signals offer better way to connect predictions to underlying economic forces, which in turn makes them more resilient in highly volatile markets.
Future work could bring in more macro-financial variables, experiment with lag structures that adapt across different market regimes. We could test whether liquidity-conditioned forecasting transfers well to other cryptocurrencies and cross-asset portfolios. This could make forecasting systems reflect how real financial decisions are made, with one eye on prices and the other on the broader economic landscape, rather than just a technical tool.